OpenStudy (anonymous):

A baseball player gets 2 hits in 9 times at bat. Estimate how many hits the player will get in the next 12 times at bat. 3 hits 5 hits 4 hits 2 hits

OpenStudy (anonymous):

I just need an answer from one of the choices that i gave...

OpenStudy (anonymous):

2/9=x/12 24=9x x=24/9, approximately 3

OpenStudy (anonymous):

ok, thanks